2x+3y=21
-3x-6y=-24

Answer by AnlytcPhil(1806)

You didn't say whether to do it by substitution or
by elimination (addition and subtraction).  I'll use
substitution

 2x + 3y =  21 
-3x - 6y = -24

Solve the second equation for x

-3x - 6y = -24
     -3x = -24 + 6y

Divide every term by -3

       x = 8 - 2y

Substitute (8-2y) for x in the OTHER (first) equation

 2x + 3y = 21

 2(8-2y) + 3y = 21

   16-4y + 3y = 21
      
       16 - y = 21
           -y = 5
            y = -5

Substitute -5 for y in x = 8-2y

            x = 8 - 2y
            x = 8 - 2(-5)
            x = 8 + 10
            x = 18 

So the solution is (x, y) = (18, -5)

Edwin
AnlytcPhil@aol.com